OpenStudy (anonymous):

Solve the equation step by step. List each step, along with the property you used during that step. –5 × 1 × 11 × 4.

OpenStudy (anonymous):

-5x1=-5 -5x11=-55 -55x4=-220

OpenStudy (anonymous):

thanks can u list the property you used

OpenStudy (anonymous):

by each step